A 2.30 kg block slides down a 16.0 m long ramp tha
Answer #1

Here,

m = 2.30 Kg

L = 16 m

theta = 55 degree

a) let the speed at the bottom is v

0.5 * m * v^2 = m * g * L * sin(theta)

0.5 * v^2 = 9.8 * 16 * sin(55)

v = 16.03 m/s

the block's speed at the bottom of the ramp is 16.03 m/s

b) let the speed of combined mass is vf

Using conseravtion of momentum

m* v = (m + 4m) * vf

vf = 16.03/5 = 3.21 m/s

the final speed is 3.21 m/s

c)

let the magnitude of force is F

Using work energy theorum

0.5 * 5m * vf^2 = F * d

0.5 * 5 * 2.3 * 3.21^2 = F * 13.3

F = 4.45 N

the magnitude of force acting is 4.45 N